In a group of 50 patients, 22 had stroke (S), 38 had concussions (C). Based on this data,

a) compute the probability of a randomly picked patient having both pain and concussions.

b) compute P(S|C)). Interpret this quantity?
Answer #1

a) n(S \cap C) = n(S) + n(C) - n(S U C)

= 22 + 38 - 50

= 60 - 50 = 10

P(S \cap C) = 10/50 = 0.2

b) P(C) = 38/50 = 0.76

P(S|C) = P(S \cap C)/P(C)

= 0.2/0.76 = 0.2632

If it is given that a randomly picked patient has concussion, then the probability that the patient had a stroke is 0.2632.